OpenStudy (anonymous):

Proof Read;Physics Solution At a point in a pipe carrying a fluid, the diameter of the pipe is 5.0 cm and the average speed of the fluid is 10 cm/s. What is the average speed, in m/s, of the fluid at a point where the diameter is 2.0 cm?

OpenStudy (anonymous):

Since I already know the diameter of the original tube which is 5.0cm, The area of the circle is calculated by (3.14)(2.5)^2=19.625cm^2 On the other hand the area of the second tube where the diameter is decreased from 5.0cm to 2.cm the area is (3.14)(1.00)^2=3.14cm^2 Therefore assuming that equation of continuity holds true in this situation 10m/s*(19.625cm^2/3.1cm^2)=10m/s*6.330=63.30m/s^2 Adjusting for sig figs of 1, Therefore velocity of the water increased from 10m/s^2 to 60m/s^2 .

OpenStudy (anonymous):

Please see if my solution is well reasoned according to the principles of continuity equation. Thanks!

OpenStudy (anonymous):

Oh by the way this is hydraulics and pneumatics

OpenStudy (anonymous):

Just so you know I am not big on those

OpenStudy (anonymous):

@nevermind_justschool

OpenStudy (anonymous):

@UsukiDoll

OpenStudy (anonymous):

In my honest opinion this line of reasoning seems pretty sound to me.

OpenStudy (nevermind_justschool):

Looks good to me :) And your reasoning sounds pretty good also. Great Job :)

OpenStudy (anonymous):

Oh yeah you think? I trust you buddy

OpenStudy (nevermind_justschool):

Yeah I think it looks pretty good :)

OpenStudy (usukidoll):

blah I'm kind of horrible at Physics :( and it's almost 3:30 am , so my mind needs to go zzz zzz

OpenStudy (anonymous):

Thanks for your try though !

OpenStudy (anonymous):

Water behaves like air does... But some properties are so different that water is used as multiplier of force.

OpenStudy (anonymous):

Did you learn about hydrostatic paradox?

OpenStudy (usukidoll):

no... I'm more geared towards the general computational math and pure math. So, anything to do with Science, I will run real fast . ^_^

OpenStudy (anonymous):

If you combine your brain and my brain it will be well balanced

OpenStudy (anonymous):

I am like screw-math-science-rocks kinda guy so